# Graph Neural Backdoor: Fundamentals, Methodologies, Applications, and Future Directions

### v1

@@ 5.3. Countermeasure Development
5.3.1. Black-box Graph Backdoor Detection
Current backdoor detection methodologies require access to training graphs and model parameters. However, to safeguard user privacy, defenders are typically unable to obtain such data. Consequently, the development of black-box detection methodology that do not rely on available samples and model parameters has become essential (Dong et al., 2021; Guo et al., 2022).
@@ 5.3.1. Black-box Graph Backdoor Detection
5.3.2. Black-box Graph Backdoor Mitigation
Based on the results of backdoor detection, users frequently petition the defense side for the removal of embedded backdoors within the model in instances where data and model parameters are unavailable (i.e., black-box conditions) for privacy protection (Wang et al., 2024d; Le Roux et al., 2024).
@@ 5.3.2. Black-box Graph Backdoor Mitigation
5.3.3. Graph Backdoor Explainability

### v2

@@ V-C Countermeasure Development
V-C1 Black-Box Graph Backdoor Detection
Current backdoor detection methodologies require access to training graphs and model parameters. However, to safeguard user privacy, defenders are typically unable to obtain such data. Consequently, the development of black-box detection methodology that do not rely on available samples and model parameters has become essential [104, 105].
@@ V-C1 Black-Box Graph Backdoor Detection
V-C2 Black-Box Graph Backdoor Mitigation
Based on the results of backdoor detection, users frequently petition the defense side for the removal of embedded backdoors within the model in instances where data and model parameters are unavailable (i.e., black-box conditions) for privacy protection [106, 107].
@@ V-C2 Black-Box Graph Backdoor Mitigation
V-C3 Graph Backdoor Explainability